𝒉(𝒙) = π₯π¨π πŸ’

(πŸπ’™

πŸ‘ βˆ’ πŸ‘)


1
Expert's answer
2021-05-07T09:45:58-0400
"h(x)=\log_4(2x^3-3)"

"h'(x)=(\log_4(2x^3-3))'=(\dfrac{\ln(2x^3-3)}{\ln4})'"

"=\dfrac{6x^2}{(2x^3-3)\ln4}"